OpenStudy (anonymous):

solve the system using the elimination method 3x+2y=-2 and 6x-y=6

OpenStudy (anonymous):

Multiply the first equation by -2. This makes it so that the x's cancel out (-6, & 6) Then solve for y.

OpenStudy (anonymous):

Did you get -6x-4y=4 ? With that first step.

OpenStudy (anonymous):

I got y=2?

OpenStudy (anonymous):

so yes then may i ask what would the next step be?

OpenStudy (anonymous):

Right! Now plug 2 back into 3x+2y=-2 for y .

OpenStudy (anonymous):

And solve for x *

OpenStudy (anonymous):

3x+4y=-2? I think I'm lost lol?!

OpenStudy (anonymous):

That is one of the original equations. You can use either, I just picked that one.

OpenStudy (anonymous):

you can plug it into 6x-y=6 instead. Both work

OpenStudy (anonymous):

ok so what would x =?

OpenStudy (anonymous):

Well, when y=2 , plug it it to solve for x. 3x+4(2)=-2 Can you solve from x from this point?

OpenStudy (anonymous):

3x+8=-2 so then you would subtract 8 to both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so therefore x=2?

OpenStudy (anonymous):

y actually equals -2 6x-y=6 -6x-4y=4 -5y=10 divide both sides by -5 y=-2 (what we said originally was wrong, y is not 2) Plug -2 back into 3x+4(-2)=-2 3x-8=-2 add 8 to both sides 3x=6 divide both sides by 3 x=2 Makes more sense. Yes x is 2

OpenStudy (anonymous):

omg yay that's how its done!!!! thanks! I got that!! :) girl, idk if this is a personal question or not but do you happen to have a facebook?

OpenStudy (anonymous):

I don't anymore, used to . But no problem :D

OpenStudy (anonymous):

well darn we could've been fb bffs lol :)